i don't understand this at all solve each of the folowing quadratic equations by completing the square.
2x^2 + 10x + 11 = 0

Answer by stanbon(75887)

2x^2 + 10x + 11 = 0
2x^2+10x = -11
2(x^2+5x+(5/2)^2) = -11 + 2*(5/2)^2
2(x+(5/2))^2 = -44/4 + 50/4
2(x+(5/2))^2 = 6/4
(x+(5/2))^2 = 6/8= 3/4
Take the square root to get:
x+(5/2) = sqrt(3/4) or x+(5/2) = -sqrt(3/4)
x = [-(5/2)+(1/2)sqrt(3)] or x = [-(5/2)-(1/2)sqrt3]
x = [-5+sqrt(3)]/2 or x=[-5-sqrt(3)]/2
================
Cheers,
Stan H.